OpenStudy (anonymous):

Two similar cylinders have a scale factor of 1.6/3. Find their volumes

OpenStudy (jdoe0001):

\(\cfrac{1.6}{3}?\)

OpenStudy (anonymous):

A) 512/3375 B) 64/225 C) 81/121 D) 64/81

OpenStudy (anonymous):

its not a decimal

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} &area&volume \\\hline\\ \textit{scale factor}=\cfrac{side1}{side2}=&\cfrac{(side1)^2}{(side2)^2}=&\cfrac{(side1)^3}{(side2)^3} \end{array}\) what do you think?

OpenStudy (anonymous):

i dont know the sides so i dont know

OpenStudy (jdoe0001):

\(\begin{array}{llll} &area&volume \\\hline\\ \textit{scale factor}=\cfrac{1.6}{3}=\cfrac{side1}{side2}=&\cfrac{(side1)^2}{(side2)^2}=&\cfrac{(side1)^3}{(side2)^3} \end{array}\) how about now?

OpenStudy (anonymous):

0.53

OpenStudy (jdoe0001):

well.. what does the fraction for the volume give you?

OpenStudy (anonymous):

is 1.6 what i plug in for side1?

OpenStudy (anonymous):

@jdoe0001 i keep getting 0.15 but thats not a fraction

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

right... wondering where the 0.15 is coming from you're not using a calculator by any chance? the result will be a fraction btw and yes, a calculator will give you the decimal form but get the fraction firstly what do you get for \(\bf \cfrac{1.6^3}{3^3}?\)